# Real-Time Video Generation with Pyramid Attention Broadcast

### v1

@@ Abstract
We present Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free approach for DiT-based video generation. Our method is founded on the observation that attention difference in the diffusion process exhibits a U-shaped pattern, indicating significant redundancy. We mitigate this by broadcasting attention outputs to subsequent steps in a pyramid style. It applies different broadcast strategies to each attention based on their variance for best efficiency. We further introduce broadcast sequence parallel for more efficient distributed inference. PAB demonstrates superior results across three models compared to baselines, achieving real-time generation for up to 720p videos. We anticipate that our simple yet effective method will serve as a robust baseline and facilitate future research and application for video generation.
Figure 1: Results and speed comparison of our method and original method. PAB can significantly boost generation speed while preserving original quality. Latency is measured on H100s. Video generation specifications: Open-Sora (2s, 480p), Open-Sora-Plan (2.7s , 512x512), Latte (2s, 512x512).
@@ 1 Introduction
Model compression methods employ techniques such as distillation (Crowley et al., 2018; Hsieh et al., 2023), pruning (Han et al., 2015; Ma et al., 2023), quantization (Banner et al., 2019; Lin et al., 2024), and novel architectures (Lin et al., 2024) to speedup deep learning models and have achieved remarkably success. Recently, they have also been proven to be effective on diffusion models (Sauer et al., 2023; Ma et al., 2024b; Chen et al., 2024a). Nevertheless, these methods usually require additional training with non-negligible computational resources and datasets, which makes model compression prohibitive and impractical especially for large-scale pre-trained models.
Most recently, people revisit the idea of cache (Smith, 1982; Goodman, 1983; Albonesi, 1999) to speedup diffusion models. Different from model compression methods, model caching methods are training-free. They alleviate redundancy by caching and reusing partial network outputs, thereby eliminating additional training. Some studies utilize high-level convolutional features for reusing purposes (Ma et al., 2024c) and efficient distributed inference (Li et al., 2024; Wang et al., 2024). Similar strategies have also been extended to specific attentions (Zhang et al., 2024; Wimbauer et al., 2024), i.e., cross attention, and standard transformers (Chen et al., 2024b).
@@
To address these challenges, we take a closer look at attentions in video DiTs and empirically obtain two observation as shown in Figure 2: (i) The attention differences between adjacent diffusion steps exhibit a U-shaped pattern, with stability in the middle 70% steps, indicating considerable redundancy for attention. (ii) Within the stable middle segment, different attentions also demonstrate varying degrees of difference. Spatial attention varies the most with high-frequency visual elements, temporal attention shows mid-frequency variations related to movements, and cross-modal attention remains the most stable, linking text with video content (Zhang et al., 2024).
Based on these observations, we propose Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free method for efficient DiT-based video generation. Our method mitigates attention redundancy by broadcasting the attention outputs to subsequent steps, thus eliminating attention computation in the diffusion process. Specifically, we apply varied broadcast ranges for different attentions in a pyramid style, based on their stability and differences as shown in Figure 2. We empirically find that such broadcast strategy can also work well to MLP layers and diffusion steps. Additionally, to enable efficient distributed inference, we propose broadcast sequence parallel, which significantly decreases generation time with much lower communication costs.
In summary, to the best of our knowledge, PAB is the first approach that achieves real-time video generation, reaching up to 20.6 FPS with a 10.5$\times$ acceleration without compromising quality. It consistently delivers excellent and stable speedup across popular open-source video DiTs, including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a). Notably, as a training-free and generalized approach, PAB has the potential to empower any future video DiTs with real-time capabilities.
@@ 2.1 Preliminaries
Figure 3: Overview of backbone of current DiT-based video generation models, which compromises Spatial Transformer block and Temporal Transformer block. Cross-attention incorporates information from the text conditioning input.
@@
Early approaches in this domain primarily leveraged GANs (Goodfellow et al., 2014), VA-VAE (Van Den Oord et al., 2017), autoregressive Transformer (Rakhimov et al., 2020) and convolution-based diffusion models (Ho et al., 2022b). However, the remarkable success of Sora (Brooks et al., 2024) has demonstrated the great potential of diffusion transformers (DiT) (Peebles & Xie, 2023) in video generation, which leads to a series of research including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), CogVideoX (Yang et al., 2024), and Latte (Ma et al., 2024a).
In this work, we focus on accelerating the DiT-based video generation models. As illustrated in Figure 3, we present the fundamental architecture of DiT-based video diffusion models. The backbone is composed of two types of Transformer blocks: spatial and temporal. Spatial Transformer blocks capture spatial information among tokens that share the same temporal index, while temporal Transformer blocks handle information across different temporal dimensions. Cross-attention enables the model to incorporate information from the conditioning input at each step, ensuring that the generated output is coherent and aligned with the given context. Note that cross-attention mechanisms are not included in the temporal blocks of some video generation models (Ma et al., 2024a).
@@ 2.2 Attention Redundancy in Video DiTs
Figure 4: a) Visualization of attention differences in Open-Sora. diff i-j represents the difference between step i and step j. b) Comparison of total attention time cost between Stable Video Diffusion (Blattmann et al., 2023) (U-Net) and Open-Sora (DiT). c) Quantitative analysis of attention differences in Open-Sora, assessed using mean squared error (MSE). The dashed line represents the average value of the corresponding attention difference.
@@
Video DiTs employ three distinct types of attention: spatial, temporal, and cross-attention. Consequently, the computational cost of attention in these models is significantly higher than in previous methods. As Figure 4(b) illustrates, the proportion of time in total for attention in video DiTs is significantly larger than in CNN approaches, which will further escalate with larger video sizes. This dramatic increase poses a significant challenge to the efficiency of video generation.
@@
To accelerate costly attention components, we conduct an in-depth analysis of their behavior. Figure 4(a) depicts the visualized differences in attention outputs across various stages. We observe that for middle segments, the differences are minimal and patterns appear similar. The first few steps show vague patterns, likely due to the initial arrangement of content. In contrast, the final steps exhibit significant differences, presumably as the model sharpens features.
@@ 2.3 Pyramid Attention Broadcast
Figure 5: Overview of Pyramid Attention Broadcast. Our method (shown on the right side) which sets different broadcast ranges for three attentions based on their differences. The smaller the variation in attention, the broader the broadcast range. During runtime, we broadcast attention results to the next several steps (shown on the left side) to avoid redundant attention computations.
Building on findings above, we propose Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free method to speedup DiT-based video generation by alleviating redundancy in attention computations. As shown in Figure 5, PAB employs a simple yet effective strategy to broadcast the attention output from some diffusion steps to their subsequent steps within the stable middle segment. Different from previous approaches that reuse attention scores (Li et al., 2023b), we choose to broadcast the entire attention module’s outputs, as we find this method to be equally effective but significantly more efficient. This approach allows us to completely bypass redundant attention computations in those subsequent steps, thereby significantly reducing computational costs. This can be formulated as:
@@
where $O_{\mathrm{attn.}}$ refers to the output of the attention module at all timesteps, $F(X_{t})$ denotes the attentions are calculated at timestep $t$ and $Y_{t}^{*}$ indicates the attentions results are broadcast from timestep $t$.
Furthermore, our research reveals that employing a single strategy across all attention types is still far from optimal speedup, as the feature and difference vary a lot for each attention as shown in Figure 2. To enhance efficiency while preserving quality, we propose to tailor broadcast ranges for each attention type, as depicted in Figure 5. The determination of the broadcast ranges is based on two key factors: the rate of change and the stability of each attention type. Attention mechanisms that exhibit more changes and fluctuations at consecutive step are assigned smaller broadcast ranges for their outputs. This adaptive methodology facilitates more efficient handling of diverse attentions within the model architecture.
@@ 2.4 Broadcast sequence parallelism
We introduce broadcast sequence parallel to enhance video generation speed based on Dynamic Sequence Parallelism (DSP) (Zhao et al., 2024), which leverages PAB’s unique characteristics to improve sequence parallelism. Sequence parallel methods distributes workload across GPUs, thus reduce generation latency. But they incur significant communication overhead for temporal attention as shown in Figure 6. By broadcasting temporal attention, we eliminate intra-module communications, substantially reducing overhead without quality loss. This enables more efficient, scalable distributed inference for real-time video generation.
@@ 3.1 Experimental setup
We select three state-of-the-art open-source DiT-based video generation models including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a) as our experimental models.
@@
We employ $\Delta$-DiT (Chen et al., 2024b) and T-GATE (Zhang et al., 2024), which are also cache-based methods as baselines in the evaluation. We show details in Appendix A.3.
@@ 3.2 Main Results
model method VBench (%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FLOPs (T) latency (s) speedup
@@
Table 1: Latency and speedup are calculated on one GPU. PABαβγ denotes broadcast ranges of spatial ($\alpha$), temporal ($\beta$), and cross ($\gamma$) attentions. Video generation specifications: Open-Sora (2s, 480p), Open-Sora-Plan (2.7s, 512x512), Latte (2s, 512x512). PSNR, SSIM, and LPIPS are calculated against the original model results. FLOPs indicate floating-point operations per video generation.
@@
Table 1 presents quality comparisons between our method and two baselines across four metrics and three models. We generate videos based on VBench’s (Huang et al., 2024) prompts, as other datasets have been shown to suffer from overfitting (Zheng et al., 2024; Lab & etc., 2024). We then evaluate VBench (Huang et al., 2024) for each method independently, and calculate other metrics including PSNR, LPIPS, and SSIM, with respect to the original model results.
Based on the results, we make the following observations: i) Our method achieves comparable or superior quality results to the two baselines while simultaneously achieving significantly higher acceleration by up to 1.58$\times$ on a single GPU. This demonstrates our method’s ability to improve efficiency with negligible quality loss. ii) Our method consistently performs well across all three models, which utilize diverse training strategies and noise schedulers, demonstrating its generalizability.
Figure 7: Speedups. We evaluate the latency and speedup achieved by PAB for single video generation across up to 8 NVIDIA H100 GPUs. The results are presented for three models utilizing broadcast sequence parallelism.
@@
Figure 7 illustrates the significant speedup achieved by our method when leveraging multiple GPUs with broadcast sequence parallelism. Our method demonstrates almost linear speedups as the GPU count increases across three different models. Notably, it achieves an impressive 10.60$\times$ speedup when utilizing 8 GPUs. These results highlight the significant reduction in communication overhead and underscore the efficacy of our broadcast sequence parallelism strategy.
@@ 3.3 Ablation Study
To thoroughly examine the characteristics of our method, we conduct extensive ablation studies. Unless otherwise stated, we apply PAB246 (the higest quality, but less speedup) to Open-Sora for generating 2s 480p videos using a single NVIDIA H100 GPU.
Table 2: Evaluation of components. w/o indicates the broadcast strategy is disabled for that component and $\Delta$ represents the corresponding increased latency.
broadcast strategy latency (s) $\Delta$ VBench (%) $\uparrow$
w/o spatial attn. 21.74 +1.87 78.45
w/o temporal attn. 23.95 +4.08 78.98
w/o cross attn. 20.98 +1.11 78.58
w/o mlp 20.27 +0.40 78.59
@@
Table 3: Broadcast object comparison. We compare the speedup, memory cost for different broadcasting object. attention outputs refer to the final output of the attention module.
broadcast object VBench (%) latency (s)
@@
As shown in Table 3.3, we compare the contribution of each component in terms of speed and quality. We disable the broadcast strategy for each component individually and measure the VBench scores and increase in latency. While the impact on VBench scores is negligible, each component contributes to the overall speedup. Spatial and temporal attentions yield the most computational savings, as they address more extensive redundancies compared to other components. Cross attention follows, offering moderate improvements despite its relatively lightweight computation. The mlp shows limited speedup due to its inherently low redundancy.
@@
i) Attention output change rates are low, as the accumulated results after attention aggregation remain similar despite pixel-level changes. This further indicates significant redundancy in attention computations. ii) Broadcasting attention scores reduces efficiency by precluding the use of optimized kernels such as FlashAttention. It also necessitates additional attention-related computations, including normalization and attention projection, which are avoided when broadcasting outputs.

### v2

@@ Abstract
We present Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free approach for DiT-based video generation. Our method is founded on the observation that attention difference in the diffusion process exhibits a U-shaped pattern, indicating significant redundancy. We mitigate this by broadcasting attention outputs to subsequent steps in a pyramid style. It applies different broadcast strategies to each attention based on their variance for best efficiency. We further introduce broadcast sequence parallel for more efficient distributed inference. PAB demonstrates up to 10.5$\times$ speedup across three models compared to baselines, achieving real-time generation for up to 720p videos. We anticipate that our simple yet effective method will serve as a robust baseline and facilitate future research and application for video generation.
Figure 1: Results and speed comparison of our and original methods. PAB can significantly boost generation speed while preserving original quality. Latency is measured on 8 H100 GPUs. Video generation specifications: Open-Sora (2s, 480p), Open-Sora-Plan (2.7s , 512x512), Latte (2s, 512x512).
@@ 1 Introduction
Model compression methods employ techniques such as distillation (Crowley et al., 2018; Hsieh et al., 2023), pruning (Han et al., 2015; Ma et al., 2023), quantization (Banner et al., 2019; Lin et al., 2024), and novel architectures (Lin et al., 2024) to speedup deep learning models and have achieved remarkably success. Recently, they have also been proven to be effective on diffusion models (Sauer et al., 2023; Ma et al., 2024b; Chen et al., 2024b). Nevertheless, these methods usually require additional training with considerable computational resources and datasets, which makes model compression prohibitive and impractical especially for large-scale pre-trained models.
Most recently, researchers revisit the idea of cache (Smith, 1982; Goodman, 1983; Albonesi, 1999) to speedup diffusion models. Different from model compression methods, model caching methods are training-free. They alleviate redundancy by caching and reusing partial network outputs, thereby eliminating additional training. Some studies utilize high-level convolutional features for reusing purposes (Ma et al., 2024c) and efficient distributed inference (Li et al., 2024; Wang et al., 2024). Similar strategies have also been extended to specific attentions (Zhang et al., 2024; Wimbauer et al., 2024), i.e., cross attention, and standard transformers (Chen et al., 2024c).
@@
To address these challenges, we take a closer look at attentions in video DiTs and empirically obtain two observations as shown in Figure 2: (i) The attention output differences between adjacent diffusion steps exhibit a U-shaped pattern, with stability in the middle 70% steps, indicating considerable redundancy for attention. (ii) Within the stable middle segment, different attention types also demonstrate various degrees of difference. Spatial attention changes the most with high-frequency visual elements, temporal attention shows mid-frequency variations related to movements, and cross-modal attention remains the most stable, linking text with video content (Zhang et al., 2024).
Based on these observations, we propose Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free method for efficient DiT-based video generation. Our method mitigates attention redundancy by broadcasting the attention outputs to subsequent steps, thus eliminating attention computation in the diffusion process. Specifically, we apply various broadcast ranges for different attentions in a pyramid style, based on their stability and differences as shown in Figure 2. We empirically find that such broadcast strategy can also work to MLP layers. Additionally, to enable efficient distributed inference, we propose broadcast sequence parallel, which significantly decreases generation time with much lower communication costs.
In summary, to the best of our knowledge, PAB is the first approach that achieves real-time video generation, reaching up to 35.6 FPS with a 10.5$\times$ acceleration without compromising quality. It consistently delivers excellent and stable speedup across popular open-source video DiTs, including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a). Notably, as a training-free and generalized approach, PAB has the potential to empower any future video DiTs with real-time capabilities.
@@ 2.1 Preliminaries
Figure 3: Overview of DiT-based video generation models, which compromises spatial and temporal transformer block. Cross attention incorporates information from text.
@@
The remarkable success of Sora (Brooks et al., 2024) has demonstrated the great potential of diffusion transformers (DiT) (Peebles & Xie, 2023) in video generation, which leads to a series of research including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a).
In this work, we focus on accelerating the DiT-based video generation models. As illustrated in Figure 3, we present the fundamental architecture of video DiTs. Different from transitional transformers, the model is composed of two types of transformer blocks: spatial and temporal. Spatial transformer blocks capture spatial information among tokens that share the same temporal index, while temporal transformer blocks handle information across different temporal dimensions. Cross-attention enables the model to incorporate information from the conditioning input at each step, ensuring that the generated output is coherent and aligned with the given context. Note that cross-attention mechanisms are not included in the temporal blocks of some video generation models (Ma et al., 2024a).
@@ 2.2 Attention Redundancy in Video DiTs
Figure 4: a) Visualization of attention differences in Latte. diff i-j represents the difference between step i and step j. b) Comparison of total attention time cost between Stable Video Diffusion (Blattmann et al., 2023) (U-Net) and Open-Sora (DiT). c) Quantitative analysis of attention differences in Open-Sora, assessed using mean squared error (MSE). The dashed line represents the average value of the corresponding attention difference.
@@
Video DiTs employ three distinct types of attentions: spatial, temporal, and cross attention. Consequently, the computational cost of attention in these models is significantly higher than in previous methods. As Figure 4(b) illustrates, the proportion of time for total attention in video DiTs is significantly larger than in CNN approaches, which will further increases with larger video sizes. This dramatic increase poses a significant challenge to the efficiency of video generation.
@@
To accelerate costly attention components, we conduct an in-depth analysis of their behavior. Figure 4(a) shows the visualized differences in attention outputs across various stages. We observe that for middle segments, the differences are minimal and patterns appear similar. The first few steps show vague patterns, likely due to the initial arrangement of content. In contrast, the final steps exhibit significant differences, presumably as the model sharpens features.
@@ 2.3 Pyramid Attention Broadcast
Figure 5: Overview of Pyramid Attention Broadcast. Our method (shown on the right side) which sets different broadcast ranges for three attentions based on their differences. The smaller the variation in attention, the longer the broadcast range. During runtime, we broadcast attention results to the next several steps (shown on the left side) to avoid redundant attention computations.
Building on the findings above, we propose Pyramid Attention Broadcast (PAB), a real-time, high quality and training-free method to speedup DiT-based video generation by alleviating redundancy in attention computations. As shown in Figure 5, PAB employs a simple yet effective strategy to broadcast the attention output from some diffusion steps to their subsequent steps within the stable middle segment of diffusion process. Different from previous approaches that reuse attention scores (Treviso et al., 2021), we choose to broadcast the entire attention module’s outputs, as we find this method to be equally effective but significantly more efficient. This approach allows us to completely bypass redundant attention computations in those subsequent steps, thereby significantly reducing computational costs. This can be formulated as:
@@
where $O_{\mathrm{attn.}}$ refers to the output of the attention module at all timesteps, $F(X_{t})$ denotes the attentions are calculated at timestep $t$ and $Y_{t}^{*}$ indicates the attentions results are broadcast from timestep $t$. We also apply similar strategy to mlp modules as depicted in Appendix A.2.2.
Furthermore, our research reveals that a single strategy across all attention types is still far from optimal, as each attention vary a lot as shown in Figure 2 and 4(c). To improve efficiency while preserving quality, we propose to tailor different broadcast ranges for each attention, as depicted in Figure 5. The determination of the broadcast ranges is based on two key factors: the rate of change and the stability of each attention type. Attention types that exhibit more changes and fluctuations at consecutive step are assigned smaller broadcast ranges for their outputs. This adaptive strategy enables more efficient handling of diverse attentions within the model architecture.
@@ 2.4 Broadcast sequence parallelism
We introduce broadcast sequence parallel, which leverages PAB’s unique characteristics to improve distributed inference speed. Sequence parallel methods (Jacobs et al., 2023; Zhao et al., 2024) distributes workload across GPUs, thus reducing generation latency. But they incur significant communication overhead for temporal attention as shown in Figure 6. By broadcasting temporal attention, we naturally eliminate extra communications, substantially reducing overhead without quality loss, which enables more efficient, scalable distributed inference for real-time video generation.
@@ 3.1 Experimental setup
We select three state-of-the-art open-source DiT-based video generation models including Open-Sora-v1.2 (Zheng et al., 2024), Open-Sora-Plan-v1.1.0 (Lab & etc., 2024), and Latte-1.0 (Ma et al., 2024a) as our experimental models.
@@
We employ $\Delta$-DiT (Chen et al., 2024c) and T-GATE (Zhang et al., 2024), which are both training-free caching methods to accelerate DiTs. We show details in Appendix A.3.
@@ 3.2 Main Results
model method VBench $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FLOPs (T) $\downarrow$ latency (s) $\downarrow$ speedup
@@
Table 1: Quality results on single GPU. PABαβγ denotes broadcast ranges of spatial ($\alpha$), temporal ($\beta$), and cross ($\gamma$) attentions. Video generation specifications: Open-Sora (2s, 480p), Open-Sora-Plan (2.7s, 512x512), Latte (2s, 512x512). PSNR, SSIM, and LPIPS are calculated against the original model results. FLOPs indicate floating-point operations per video generation.
@@
Table 1 presents quality comparisons between our method and baselines across four metrics and three models. We generate videos based on VBench’s (Huang et al., 2024) prompts. Then evaluate VBench for each method, and calculate other metrics including PSNR, LPIPS, and SSIM with respect to the original results. PABαβγ denotes broadcast ranges of spatial ($\alpha$), temporal ($\beta$), and cross ($\gamma$) attentions. More experiments on other datasets can be found in Appendix B.1.
Based on the results, we make the following observations: i) Our method achieves superior quality results compared with two baselines while simultaneously achieving significantly higher acceleration by up to 1.58$\times$ on a single GPU. This demonstrates our method’s ability to improve efficiency with negligible quality loss. ii) Our method consistently performs well across all three models, which utilize diverse training strategies and noise schedulers, demonstrating its generalizability.
Figure 7: Speedups. We evaluate the latency and speedup achieved by PAB246/PAB235 (the strategy with best quality, but less speedup) for single video generation across up to 8 NVIDIA H100 GPUs. The results are presented for three models utilizing broadcast sequence parallelism. The multiple GPUs’ speedup is compared with single GPU’s speed.
@@
Figure 7 illustrates the significant speedup achieved by our method when leveraging multiple GPUs with broadcast sequence parallelism. Our method demonstrates almost linear speedups as the GPU number increases across three different models. Notably, it achieves an impressive 10.50$\times$ speedup when utilizing 8 GPUs. These results highlight the significant reduction in communication overhead and underscore the efficacy of our broadcast sequence parallelism strategy.
@@ 3.3 Ablation Study
To thoroughly examine the characteristics of our method, we conduct extensive ablation studies. Unless otherwise stated, we apply PAB246 (the best quality, but less speedup) to Open-Sora for generating 2s 480p videos using a single NVIDIA H100 GPU.
Table 2: Evaluation of components. w/o indicates that the broadcast strategy is disabled only for the corresponding component.
broadcast strategy latency (s) $\downarrow$ VBench $\uparrow$
[…]
@@
Table 3: Broadcast object comparison. We compare the speedup and effect for attention map and attention outputs. attention outputs refer to the final output of attention module. attention scores denotes attention score map.
broadcast object VBench $\uparrow$ latency (s) $\downarrow$
@@
As shown in Table 3.3, we compare the contribution of each component in terms of speed and quality. We disable the broadcast strategy for each component individually and measure the VBench scores and increase in latency. While the impacts on VBench scores are negligible, all components contribute to the overall speedup. Spatial and temporal attentions yield the most computational savings, as they address more extensive redundancies compared to other components. Cross attention follows, offering moderate improvements despite its relatively lightweight computation. The mlp shows limited speedup due to its inherently low redundancy.
@@
i) Attention output change rates are low, as the accumulated results after attention aggregation remain similar despite pixel-level changes. This further indicates significant redundancy in attention computations. ii) Broadcasting attention scores prevents the use of efficient attention kernels such as FlashAttention (Dao et al., 2022). It also requires complete attention-related computations, including attention calculation and linear projection, which are avoided when broadcasting outputs.